OpenStudy (anonymous):

Which of the following is equivalent to log8 16 note: the 8 is really small.

OpenStudy (anonymous):

a: 2 b: 1.248 c: 4/3 d:3/4

OpenStudy (ash2326):

You mean \[\log_8 16\]?

OpenStudy (anonymous):

yes sir

OpenStudy (ash2326):

We have \[\log_8 16\] it means we need to find x for which \[8^x=16\] Can you change the base to 2?

OpenStudy (anonymous):

Which of the following is equivalent to log 8 16 thats all it says

OpenStudy (ash2326):

That's fine but do you know the properties of logarithms?

OpenStudy (anonymous):

i have no clue what a logarithm is.

OpenStudy (ash2326):

If we have \[\log_x y=c\] then this implies \[y=x^c\] Do you get this?

OpenStudy (anonymous):

i don't understand that.

OpenStudy (anonymous):

i need to finish like 80 problem so i pass my math class and i don't understand them

OpenStudy (anonymous):

problems

OpenStudy (anonymous):

is the answer a?

OpenStudy (ash2326):

@zackwashere I gotta go now, Sorry I'll ask someone to help you! @lgbasallote would you help here?

OpenStudy (anonymous):

later

OpenStudy (anonymous):

Howdy, Zack =)

OpenStudy (anonymous):

yello

OpenStudy (anonymous):

Let me show you an example. \[\large \log_2{4}\] Means, "What power can I raise 2 to to get 4?" So we need to come up with a number that we can raise 2 to to get 4, and we try 2. \[2^2 = 4\] So that means that 2 works, and we get \[\large \log_2{4}=2\]

OpenStudy (anonymous):

Here's another example. \[\large \log_2{8}\] means that we need some power of 2 that will give us 8. \[2^3 = 8\] so \[\large \log_2{8}=3\]

OpenStudy (anonymous):

i think the answer is a am i right?

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

log_a(b) = logb/loga

OpenStudy (zzr0ck3r):

then you can change the base

OpenStudy (zzr0ck3r):

because you now have a ratio, so a base like 2 would be nice

OpenStudy (anonymous):

If you choose a, then you are saying that \[\large \log_8{16} = 2\] Which would mean that \[\large 8^2 = 16\] Is that true? No, because 8^2 = 64

OpenStudy (anonymous):

i have no clue what you guys are talking about

OpenStudy (zzr0ck3r):

is this for a class? have you studied? because there are lots of log rules and it will take some time to get used to them, but you have to practice.

OpenStudy (anonymous):

i fail at math and i need to finish a few math problems to pass the class

OpenStudy (anonymous):

my friend said this site gave people answers

OpenStudy (anonymous):

Naw. That ain't me.

OpenStudy (anonymous):

well i know it's not a >.>

OpenStudy (zzr0ck3r):

If you jsut want the answer type this into wolframalpha.com log_8(16)

OpenStudy (zzr0ck3r):

but thats not going to help later in life:)

OpenStudy (anonymous):

gnarly man!

OpenStudy (anonymous):

you just saved me a crap load of time

OpenStudy (anonymous):

4/3!